You were hired as a consultant to Okland Company, whose target capital structure is 51% debt, and 49% common equity. The interest rate on new debt is 5.40%, the cost of retained earnings is 13.60%, and the tax rate is 22% The firm will not be issuing any new stock. What is Okland's WACC?
Answer #1

After-tax cost of debt=5.4*(1-tax rate)

=5.4*(1-0.22)=4.212%

WACC=Respective costs*Respective weight

=(4.212*0.51)+(0.49*13.6)

=8.81212%